uppose that a linear system of equations in unknowns x, y, and z has the following augmented matrix.
1 -1 2 -2
-4 -2 3 1
3 -2 4 -3


Use Gauss-Jordan elimination to solve the system for x, y, and z.

Problem #7:

Enter the values of x,y, and z here,
in that order, separated by commas.
Answer #1

The software output for this problem is:

Result: (x - y +2 2 = -2,-4x - 2y +3 z = 1, 3x - 2y + 4z = -3) Enlarge Data Customize A Plain Text Solution: Step-by-step sol

Hence,

Answer with values of x, y and z will be:

1, -19, -11